OpenStudy (anonymous):

From 7 boys and 4 girls, how many different committees can be selected consisting of 3 boys and 2 girls? please help!!!!

OpenStudy (anonymous):

C(7,3) * C(4,2) 35 * 6 210

OpenStudy (anonymous):

Boys: 3/7 and girls: 2/4= 1/2

OpenStudy (anonymous):

@sean_davis0821 She's not looking for a probability. She has to count the number of possible committees.

OpenStudy (anonymous):

yee thank uuu :))) I did 35+6 =41 this way, so we should multiple them :))) thanks so much!!

OpenStudy (anonymous):

No problem. Yes...multiply them...just like if you had 4 pants and 3 shirts, how many outfits could you put together? Not 4 + 3....It's 4 times 3 = 12 possible combinations.

OpenStudy (anonymous):

like accounting :))